# How Far Are We from Intelligent Visual Deductive Reasoning?

### v2

@@ Abstract
Vision-Language Models (VLMs) such as GPT-4V have recently demonstrated incredible strides on diverse vision language tasks. We dig into vision-based deductive reasoning, a more sophisticated but less explored realm, and find previously unexposed blindspots in the current SOTA VLMs. Specifically, we leverage Raven’s Progressive Matrices (RPMs), to assess VLMs’ abilities to perform multi-hop relational and deductive reasoning relying solely on visual clues. We perform comprehensive evaluations of several popular VLMs employing standard strategies such as in-context learning, self-consistency, and Chain-of-thoughts (CoT) on three diverse datasets, including the Mensa IQ test, IntelligenceTest, and RAVEN. The results reveal that despite the impressive capabilities of LLMs in text-based reasoning, we are still far from achieving comparable proficiency in visual deductive reasoning. We found that certain standard strategies that are effective when applied to LLMs do not seamlessly translate to the challenges presented by visual reasoning tasks. Moreover, a detailed analysis reveals that VLMs struggle to solve these tasks mainly because they are unable to perceive and comprehend multiple, confounding abstract patterns in RPM examples.
@@ 1 Introduction
  1. 1.
    We set up a framework for systematically evaluating VLMs on RPM problems. We evaluated several SOTA open-source and closed-source VLMs on three diverse datasets, including the Mensa IQ test, IntelligenceTest, and RAVEN, providing a comprehensive assessment of their performance. The results indicate that although LLMs exhibit impressive capabilities in text-based reasoning, such proficiency has not been achieved in image-based reasoning. The code and evaluation datasets have been released to facilitate future investigation and improvement over VLMs.  [^1] Our code is available at https://github.com/apple/ml-rpm-bench
  2. 2.
@@ 2 Related Work
Many text-based reasoning tasks and benchmarks have been introduced to evaluate LLMs in various domains (Huang & Chang, 2022) such as general knowledge (Hendrycks et al., 2020), math reasoning (Cobbe et al., 2021), commonsense reasoning (Geva et al., 2021; Clark et al., 2018), factual reasoning (Laban et al., 2023), and coding (Chen et al., 2021). Some noteworthy examples of these works are BIG-bench (Srivastava et al., 2022), HELM (Liang et al., 2022), SuperGLUE (Sarlin et al., 2020), and LAMA (Petroni et al., 2019).
Visual reasoning evaluation
@@
Previous work on visual reasoning tasks has primarily focused on tasks such as visual question answering (VQA) Antol et al. (2015) and image captioning. These tasks involve answering questions about images or generating natural language descriptions of visual content. Researchers have also examined the ability of models to understand the relational and compositional aspects of objects in images. Datasets like CLEVR (Johnson et al., 2017) and SHAPES (Andreas et al., 2016) assess visual reasoning abilities such as counting, comparing, logical reasoning, and storing information in memory. As the VLMs abilities to perform visual reasoning have evolved so have the benchmarks. New benchmarks, like MMMU (Yue et al., 2023) and MathVista (Lu et al., 2023) have been developed that test the models’ ability to emulate human-like understanding of scenes and objects in images and videos. These benchmarks include areas such as scene text understanding (Sidorov et al., 2020), formulation (Lu et al., 2024), table and chart interpretation (Lu et al., 2024), the comprehension of visual stimuli (Yang et al., 2023), geometric reasoning (Ahrabian et al., 2024), spatial reasoning (Chen et al., 2024), and facial expression comprehension and reasoning (Yang et al., 2023).
Deductive reasoning
@@
Deductive reasoning evaluation and benchmarks have been conducted for both textual and visual domains. Two notable examples are GuessWhat?! (De Vries et al., 2017) and ReferIt (Kazemzadeh et al., 2014), which assess the visual reasoning abilities of the models being tested. More recently, LMRL Gym (Abdulhai et al., 2023) and Entity Deduction Arena (Zhang et al., 2023) have been introduced as methods to evaluate the ability of LLMs to perform multi-turn deductive reasoning tasks. Another relevant task is ARC (Acquaviva et al., 2022) which shares similarities with RPMs, as they both require correctly inferring unseen outputs based on given examples. Comparing with ARC, RPMs are abstract and requires intricate analogical and relational reasoning.
@@ 3 Experiment Setting
3.1 Dataset
In our paper, we employed three RPMs datasets. The Mensa Test [^2] https://www.mensa.org/public/mensa-iq-challenge consists of 35 questions with progressive levels of difficulty. For the purpose of 1-shot learning, we used the first question as an in-context example and reserved the remaining 34 questions for evaluation. The IntelligenceTest (IT) [^3] https://www.intelligencetest.com/questions provided an IQ test encompassing verbal, pattern recognition, math, and structural components. We specifically focused on pattern recognition, which solely comprised RPMs problems and included 66 examples. Additionally, we incorporated the RAVEN dataset (Zhang et al., 2019) for evaluation. The RAVEN dataset employs a generative model to create RPMs problems using a hierarchical pipeline. The test dataset of RAVEN contains 14,000 examples, covering 7 types of distinct figural configurations that incorporate different layouts, shapes, and relational structures. In this work, we generate 140 new samples, 20 samples for each figural configuration.
3.2 Models
We compared various VLMs that represent the state-of-the-art for both closed-source and open-source models, including GPT4-V (gpt-4-vision-preview) (OpenAI, 2023), Gemini-pro (Team et al., 2023), Qwen-VL-Max (Bai et al., 2023) and LLaVA-1.5-13B (Liu et al., 2023). We use the default sampling method for each of the tested VLMs in our generation process, show in Appendix A.2.
3.3 Prompts
We prompt the model with the instruction followed by the query image. For the Mensa set we used the following prompt:
You can see a grid of 9 boxed, one of which is empty (marked as ?). You have to choose which of the 6 alternative shapes (A-F) should be placed in the empty box in order to complete the pattern that connects the shapes. Finally, provide your prediction as Answer: “X”. {query image}
Appendix A.1 contains the prompts used for all three datasets for different settings. In Section 5.4, we explain why we chose the particular order of the prompt and query image order.
@@ 4.1 Evaluation of VLMs on Visual Deductive Reasoning
We analyze the entropy of model predictions in order to assess the uncertainty inherent in their predictive distribution. For the choices set $C$, the Entropy is defined as $S=-\sum_{i\in C}p_{i}\log p_{i}$. If the model consistently predicts a single answer, it is has an entropy of 0. If it randomly guesses, the entropy reaches the upper bound shown in the Table 1. We see that GPT-4 and Gemini Pro exhibit a greater diversity of answers, which is also reflected in the greater diversity in recognizing and attempting to identify various patterns. On the other hand, LLaVA and QWen-VL produce more deterministic predictions, resulting in lower entropy.
We observed that all the VLMs tend to be confident while being wrong. Interestingly, we observed that even when the entropy was high, models tried to provide a nonsensical rational, instead of acknowledging their inability to perform the task; this was observed to happen more often with models that had higher entropy. All the tested models never express any level of uncertainty by using words like “likely” or “maybe”. This excessive confidence can presumably be attributed to the model pretraining and instruction finetuning steps, which typically do not involve calibrating the model for uncertainty. Instead, the models are encouraged to generate uncertain content, leading to more errors in aggregating in the generated output.
4.2 Do some standard strategies used in LLMs translate effectively to visual deductive reasoning?
We tried two strategies that have been shown to be effective in various tasks related to text LLMs: 1) One-shot (1-shot) (Brown et al., 2020) involves providing a single in-context RPM example and its solution to the VLMs. 2) Self-consistency (SC) (Wang et al., 2022) entails sampling multiple responses and selecting the answer that receives the majority vote. The results are shown in Table 2.
@@ 4.2 Do some standard strategies used in LLMs translate effectively to visual deductive reasoning?
The performance of the 1-shot evaluation did not demonstrate a significant improvement compared to the 0-shot evaluation. Specifically, we observed only a marginal 1% enhancement for the IntelligenceTest dataset, while encountering a decrease of 2-4% in accuracy for the Mensa test. Surprisingly, all the tested models, including GPT-4V and Gemini, struggle with a high failure rate even when the in-context example is identical to the current task being solved. This is peculiar because powerful LLMs usually exhibit the ability to analogize and copy the in-context example when provided with the same query. We observed accuracy ranging from 10% to 20% for these in-context examples across different datasets, which is comparable to the accuracy when a different example is used as the in-context example.
@@ 5 What limits the performance of the VLMs?
Figure 2: Three manually created RPM problems evaluated for text description augmentation, illustrating varying levels of difficulty: easy, medium, and hard. The correct answers are “F, F, F”.
@@
We investigate why VLMs fail to reach human-level performance in answering even simple questions that are intuitive to humans. For this purpose, as a case study, we manually created three RPMs with varying degrees of difficulty, as depicted in Figure 2. To conduct a fine-grained analysis and diagnosis of the VLM’s inability to perform this task of visual deductive reasoning with RPMs, we decompose the evaluation into three consecutive stages:
  • Perception: assess if the model can understand and describe the patterns in the RPMs.
  • Deductive reasoning: evaluate if the model can discern and articulate underlying rules.
  • Hypothesis verification: examine the model’s proficiency in formulating a plausible hypothesis for the missing pattern and identifying a matching option among alternatives.
@@ 5.1 How good is the VLM’s perception on this task?
We first asked the model to describe the RPM figures, to assess if they understood the images that were provided as part of the problem. Surprisingly, even though VLMs are astoundingly accurate in describing commonplace images, they seemed to be quite unsuccessful at accurately describing even the simpler abstract patterns we gave them. The generated descriptions contained numerous errors across all the tested models, as exemplified by results from GPT-4V in Table 4.
We identified two major issues for this blind spot of VLMs:
  1. 1.
@@
These two issues are prevalent across all the tested methods and dataset. When the patterns contain multiple elements and are more detailed, these issues become more severe.
Can decomposing the RPMs into each single pattern from the grid enhance perception?
Presumably, by decomposing the patterns into individual components, we can eliminate the confounding errors. To investigate this, we first segmented each of the three manual examples shown in Figure 2, into 9 individual question patterns and 6 candidate patterns. We then used a new prompt A.1 for GPT-4V to read both the full image and the segmented patterns to infer the answer. In this way, we found GPT-4V would describe each pattern more accurately. The descriptions of the M-Medium RPM can be found in Table 4. We conducted 10 tests for each RPM and report the accuracy comparison with and without segmentation in Figure 3. We also verify the segmentation impact using the Raven dataset (140 examples). We got 16.4% accuracy for segmented RPMs and 11.4% for non-segmented RPMs. The results demonstrate a significant reduction in confounding errors, confirming the issues discussed earlier.
Hallucination
We have also observed the model generating hallucinations in the descriptions, particularly when it comes to counting. For instance, in Table 4 (M-Medium, top right), the model erroneously states that there are four dots when, in reality, there are only three.
Data distribution aspect
VLMs are presumably trained primarily on real world images, which may cause them to be less sensitive to abstract patterns. We did not try to finetune the model to test this hypothesis because of the limited availability of diverse RPM data. We believe that additional finetuning could potentially improve the performance. However, we hypothesize that finetuning the model with RPMs might not entirely eliminate the compounding and confounding errors, as they appear to be inherent limitations of the VLMs from training.
@@ 5.2 How good is the VLM’s deductive reasoning on this task?
Table 5: Accuracy and breakdown of GPT-4V variants with augmented text description across different RPMs. Each combination is ran for 10 repetitions. The correct answer “F” is marked in color.
@@
Next, we examine whether textual information alone is sufficient by removing the visual information. The results, shown in Table 5 (GPT-4V + Oracle Desc. - Visual), are intriguing. Without visual information, the models can maintain a similar level of performance for M-Easy and M-Medium RPMs. Notably the result solely rely on the textual information of the input is superior to the GPT-4V baseline, which mostly rely on visual information of the input. However, as the tasks become more challenging (M-Hard RPM), the models start to struggle. The performance is also worse than GPT-4V baseline. This suggests that for tasks that involve complex spatial layouts and relational reasoning, text alone may be insufficient and potentially confusing, while visual cues may provide additional visual alignment and better comparative attention. In such cases, visual information and textual clues would complement each other and work in synergy to achieve the optimal performance. Interestingly, when we provide GPT-4V with an incorrect description, there is around an 80% chance that the model recognizes the mismatch between the text and the image and responds with statements such as: “I believe there has been a misinterpretation of the provided image”. The model, nevertheless, still generates some rationale which seems adhere more closely to the text description than to the visual cues.

### v3

@@ Abstract
Vision-Language Models (VLMs) have recently demonstrated incredible strides on diverse vision language tasks. We dig into vision-based deductive reasoning, a more sophisticated but less explored realm, and find previously unexposed blindspots in the current SOTA VLMs. Specifically, we leverage Raven’s Progressive Matrices (RPMs), to assess VLMs’ abilities to perform multi-hop relational and deductive reasoning relying solely on visual clues. We perform comprehensive evaluations of several popular VLMs employing standard strategies such as in-context learning, self-consistency, and Chain-of-thoughts (CoT) on three diverse datasets, including the Mensa IQ test, IntelligenceTest, and RAVEN. The results reveal that despite the impressive capabilities of LLMs in text-based reasoning, we are still far from achieving comparable proficiency in visual deductive reasoning. We found that certain standard strategies that are effective when applied to LLMs do not seamlessly translate to the challenges presented by visual reasoning tasks. A detailed analysis reveals that VLMs struggle to solve these tasks mainly because they are unable to perceive and comprehend multiple, confounding abstract patterns in RPM examples.
@@ 1 Introduction
  • We set up a framework for systematically evaluating VLMs on RPM problems. We evaluated several SOTA open-source and closed-source VLMs on three diverse datasets, including the Mensa IQ test, IntelligenceTest, and RAVEN, providing a comprehensive assessment of their performance. The results indicate that although LLMs exhibit impressive capabilities in text-based reasoning, such proficiency has not been achieved in image-based reasoning. The code and evaluation datasets have been released to facilitate future investigation and improvement over VLMs.
@@ 2 Related Work
General LLM Reasoning benchmarks Many text-based reasoning tasks and benchmarks have been introduced to evaluate LLMs in various domains (Huang & Chang, 2022) such as general knowledge (Hendrycks et al., 2020), math reasoning (Cobbe et al., 2021), commonsense reasoning (Geva et al., 2021; Clark et al., 2018), factual reasoning (Laban et al., 2023), and coding (Chen et al., 2021). Some noteworthy examples of these works are BIG-bench (Srivastava et al., 2022), HELM (Liang et al., 2022) and SuperGLUE (Sarlin et al., 2020).
@@
Visual reasoning evaluation Previous work on visual reasoning tasks has primarily focused on tasks such as visual question answering (VQA) Antol et al. (2015) and image captioning. These tasks involve answering questions about images or generating natural language descriptions of visual content. Researchers have also examined the ability of models to understand the relational and compositional aspects of objects in images. Datasets like CLEVR (Johnson et al., 2017) and SHAPES (Andreas et al., 2016) assess visual reasoning abilities such as counting, comparing, logical reasoning, and storing information in memory. As the VLMs abilities to perform visual reasoning have evolved so have the benchmarks. New benchmarks, like MMMU (Yue et al., 2023) and MathVista (Lu et al., 2023) have been developed that test the models’ ability to emulate human-like understanding of scenes and objects in images and videos. These benchmarks include areas such as scene text understanding (Sidorov et al., 2020; Schiappa et al., 2024), formulation (lu2024mathvista), table and chart interpretation (lu2024mathvista), the comprehension of visual stimuli (Yang et al., 2023), geometric reasoning (Ahrabian et al., 2024), spatial reasoning (Chen et al., 2024), and facial expression comprehension and reasoning (Yang et al., 2023).
This paper focuses on RPMs. Our goal was to simulate a more holistic challenge that mirrors scenarios where generalist VLMs must navigate through even unseen and unfamiliar scenarios. Although the RPMs were designed for humans, they represent a fundamental type of visual-spatial reasoning that artificial intelligence systems, particularly those aimed at achieving general intelligence, should also be able to perform well on.
@@
Deductive reasoning Deductive reasoning evaluation and benchmarks have been conducted for both textual and visual domains. Two notable examples are GuessWhat?! (De Vries et al., 2017) and ReferIt (Kazemzadeh et al., 2014), which assess the visual reasoning abilities of the models being tested. More recently, LMRL Gym (Abdulhai et al., 2023) and Entity Deduction Arena (Zhang et al., 2023) have been introduced as methods to evaluate the ability of LLMs to perform multi-turn deductive reasoning tasks. Another relevant task is ARC (Acquaviva et al., 2022) which shares similarities with RPMs, as they both require correctly inferring unseen outputs based on given examples. Comparing with ARC, RPMs are abstract and requires intricate analogical and relational reasoning.
Using RPMs to assess VLM is not rare. For instance, many general VLM benchmarks like MATHVISTA (Lu et al., 2023) and MMMU (Yue et al., 2023) uses RPMs problems to probe models’ reasoning and understanding ability of complex patterns. When GPT-4v came out, technical evaluation work (Yang et al., 2023) also probed its performance on abstract patterns, especially the RPMs. Different from them, our work dives deep into scrutinizing the blindspot, and the underlying issues when VLMs deals with these abstract patterns, showing some potential directions for future improvement.
@@ 3 Experiment Setting
Dataset In our paper, we employed three RPMs datasets. The Mensa Test [^1] https://www.mensa.org/public/mensa-iq-challenge consists of 35 questions with progressive levels of difficulty. For the purpose of 1-shot learning, we used the first question as an in-context example and reserved the remaining 34 questions for evaluation. The IntelligenceTest (IT) [^2] https://www.intelligencetest.com/questions provided an IQ test encompassing verbal, pattern recognition, math, and structural components. We specifically focused on pattern recognition, which solely comprised RPMs problems and included 66 examples. Additionally, we incorporated the RAVEN dataset (Zhang et al., 2019) for evaluation. The RAVEN dataset employs a generative model to create RPMs problems using a hierarchical pipeline. The test dataset of RAVEN contains 14,000 examples, covering 7 types of distinct figural configurations that incorporate different layouts, shapes, and relational structures. In this work, we generate 140 new samples, 20 samples for each figural configuration.
Models We compared various VLMs that represent the state-of-the-art for both closed-source and open-source models, including GPT4-V (gpt-4-vision-preview) (OpenAI, 2023), Gemini-pro (Team et al., 2023), Qwen-VL-Max (Bai et al., 2023), LLaVA-1.5-13B and LLaVA-1.6-34B(Liu et al., 2023). We use the default sampling method for each of the tested VLMs in our generation process, show in Appendix A.3.
Prompts We prompt the model with the instruction followed by the query image. We provide the prompt in the Appendix A.2.
@@ 4.1 Evaluation of VLMs on Visual Deductive Reasoning
We analyze the entropy of model predictions in order to assess the uncertainty inherent in their predictive distribution. For the choices set $C$, the Entropy is defined as $S=-\sum_{i\in C}p_{i}\log p_{i}$. If the model consistently predicts a single answer, it is has an entropy of 0. If it randomly guesses, the entropy reaches the upper bound shown in the Table 1. For Self-consistency model, we bootstrapped 1000 leave-one-out repetitions from 5 answers per question and calculated entropy over aggregated predictions from each repetition.
We see that GPT-4 and Gemini Pro exhibit a greater diversity of answers, which is also reflected in the greater diversity in recognizing and attempting to identify various patterns. On the other hand, LLaVA and QWen-VL produce more deterministic predictions, resulting in lower entropy.
Interestingly, we observed that even when the entropy was high, models tried to provide a nonsensical rationale, instead of acknowledging their inability to perform the task; this was observed to happen more often with models that had higher entropy. All the tested models never express any level of uncertainty by using words like “likely” or “maybe”. This excessive confidence can presumably be attributed to the model pretraining and instruction finetuning steps, which typically do not involve calibrating the model for uncertainty. Instead, the models are encouraged to generate uncertain content, leading to more errors in aggregating in the generated output.
4.2 Do standard strategies in LLMs translate effectively to visual deductive reasoning?
We tried two strategies effective in LLMs: 1) 1-shot (Brown et al., 2020) prompts in-context RPM example and its solution to the VLMs. 2) Self-consistency (SC) (Wang et al., 2022) samples multiple responses and selecting the majority voted answer.
@@ 4.2 Do standard strategies in LLMs translate effectively to visual deductive reasoning?
The performance of the 1-shot evaluation, shown in Table 2, did not demonstrate improvement compared to the 0-shot evaluation. Specifically, we observed only a marginal 1% enhancement for the IntelligenceTest dataset, while encountering a decrease of 2-4% in accuracy for the Mensa test. Surprisingly, all the tested models, including GPT-4V and Gemini, struggle with a high failure rate even when the in-context example is identical to the current task being solved. This is peculiar because powerful LLMs usually exhibit the ability to analogize and copy the in-context example when provided with the same query. We observed accuracy ranging from 10% to 20% for these in-context examples across different datasets, which is comparable to the accuracy when a different example is used as the in-context example.
@@ 5 What limits the performance of the VLMs?
Figure 2: Three manually created RPM problems evaluated for text description augmentation, illustrating varying levels of difficulty. The correct answers are “F, F, F”.
@@
We investigate why VLMs fail to reach human-level performance in answering even simple questions that are intuitive to humans. For this purpose, as a case study, we manually created three RPMs with varying degrees of difficulty, as depicted in Figure 2. The manually curated examples are similar to instances in the larger datasets (e.g. Mensa), whose example cannot be presented due to copyright issues. We aim to use these examples as qualitative probes to showcase the blind spots in depth, and show that the issues are ubiquitous in VLMs, spanning stages of perception, reasoning, and verification. To conduct a fine-grained analysis and diagnosis of the VLM’s inability to perform this task of visual deductive reasoning with RPMs, we decompose the evaluation into three consecutive stages: 1) Perception: assess if the model can understand and describe the patterns in the RPMs; 2) Deductive reasoning: evaluate if the model can discern and articulate underlying rules; 3) Hypothesis verification: examine the model’s proficiency in formulating a plausible hypothesis for the missing pattern and identifying a matching option among alternatives.
@@ 5.1 How good is the VLM’s perception on this task?
We first asked the model to describe the RPM figures, to assess if they understood the images that were provided as part of the problem. Surprisingly, even though VLMs are astoundingly accurate in describing commonplace images, they seemed to be quite unsuccessful at accurately describing even the simpler abstract patterns we gave them. The generated descriptions contained numerous errors across all the tested models, as exemplified by results from GPT-4V in Table 4. More examples are shown in Appendix A.5. We identified two major issues for this blind spot of VLMs:
@@
These two issues are prevalent across all the methods and datasets. When the patterns contain multiple elements and are more detailed, these issues become severer.
Can decomposing the RPMs into each single pattern from the grid enhance perception? Presumably, by decomposing the patterns into individual components, we can eliminate the confounding errors. To investigate this, we first segmented each of the three manual examples shown in Figure 2, into 9 individual question patterns and 6 candidate patterns. We then used a new prompt A.2 for GPT-4V to read both the full image and the segmented patterns to infer the answer. In this way, we found GPT-4V would describe each pattern more accurately. The descriptions of the M-Medium RPM can be found in Table 4. We conducted 10 tests for each RPM and report the accuracy comparison with and without segmentation in Figure 3. We also verify the segmentation impact using the Raven dataset (60 examples). We got 21.2% accuracy for segmented RPMs and 7.4% for non-segmented RPMs. The results demonstrate a significant reduction in confounding errors, confirming the issues discussed earlier.
Hallucination We have also observed the model generating hallucinations in the descriptions, particularly regarding counting. For instance, in Table 4 (M-Medium, top right), the model erroneously states that there are four dots when, in reality, there are only three.
Data distribution aspect VLMs are presumably trained primarily on naturalist images, which may cause them to be less sensitive to abstract patterns. However, evaluation tasks on foundation models often involve significant distribution shifts or completely novel tasks, and foundational models like VLMs are expected to have some capacity for task generalization and zero-shot capabilities. RPM puzzles challenge VLMs to engage in abstract reasoning and pattern recognition, serving as a controlled test for perception, self-reflection, and deductive reasoning abilities. Additionally, IQ evaluations for humans already use abstract visual puzzles, establishing a comparative connection to human performance.
While we believe that additional finetuning could potentially improve the performance, we hypothesize that finetuning the model with RPMs might not entirely eliminate the compounding and confounding errors, as they appear to be inherent limitations of the VLMs from training.
@@ 5.2 How good is the VLM’s deductive reasoning on this task?
Table 5: Breakdown of GPT-4V variants with augmented text description across different RPMs. Each combination is ran for 10 repetitions. The correct answer “F” is marked in color. Given budget constraint we only test on the 3 manual examples for this analysis.
@@
Next, we examine whether textual information alone is sufficient by removing the visual information. The results, shown in Table 5 (GPT-4V + Oracle Desc. - Visual), are intriguing. Without visual information, the models can maintain a similar level of performance for M-Easy and M-Medium RPMs. Notably the result solely rely on the textual information of the input is superior to the GPT-4V baseline, which mostly rely on visual information of the input. However, as the tasks become more challenging (M-Hard RPM), the models start to struggle. The performance is also worse than GPT-4V baseline. This suggests that for tasks that involve complex spatial layouts and relational reasoning, text alone may be insufficient and potentially confusing, while visual cues may provide additional visual alignment and better comparative attention. In such cases, visual information and textual clues would complement each other and work in synergy to achieve the optimal performance. Interestingly, when we provide GPT-4V with an incorrect description, there is around an 80% chance that the model recognizes the mismatch between the text and the image and responds as: “There has been a misinterpretation of the provided image”. The model, nevertheless, still generates some rationale which seems adhere more closely to the text description than to the visual cues.